OpenStudy (bluebeta):

Who has time to look over 10 questions for me? They're all answered, I just need to know if I'm correct. Please and Thank you.

OpenStudy (bluebeta):

It'll take me a second upload them.

OpenStudy (anonymous):

ill try my best what grade level

OpenStudy (bluebeta):

12th

OpenStudy (anonymous):

never mind lol onlt a sophmore

OpenStudy (anonymous):

so sorry

OpenStudy (anonymous):

i can look them up if ud like if noone else can help you

OpenStudy (bluebeta):

I've already tried to look some up. I'm gonna be posting the work in a second, thanks though.

OpenStudy (bluebeta):

Can you try to find this one!? Write the equation of the conic section with the given properties: A hyperbola with vertices (0,-6) and (0,6) and asymptotes y=3/4x and y=-3/4x .

OpenStudy (bluebeta):

There's the 10 questions

OpenStudy (anonymous):

k ill look over them one minute

OpenStudy (bluebeta):

ty :)

OpenStudy (bluebeta):

@dolphinlover Are you viewing too?

OpenStudy (bluebeta):

@RCCB

OpenStudy (anonymous):

im in 7th srry

OpenStudy (anonymous):

@ganeshie8

OpenStudy (bluebeta):

oh ok

OpenStudy (bluebeta):

@AlexandervonHumboldt2

OpenStudy (anonymous):

they look good i justa had my brother check them hes in twelth and he said that it looks perfect

OpenStudy (anonymous):

good job

ganeshie8 (ganeshie8):

heyy

OpenStudy (bluebeta):

@ganeshie8 hey hey could you check over too?

OpenStudy (bluebeta):

@i_suck_at_school thanks!!

OpenStudy (bluebeta):

I need #9 though :(

ganeshie8 (ganeshie8):

Can you try to find this one!? Write the equation of the conic section with the given properties: A hyperbola with vertices (0,-6) and (0,6) and asymptotes y=3/4x and y=-3/4x .

ganeshie8 (ganeshie8):

that one ?

OpenStudy (bluebeta):

Yess

ganeshie8 (ganeshie8):

try \[\dfrac{y^2}{36}-\dfrac{x^2}{64}=1\]

ganeshie8 (ganeshie8):

Notice that when \(x=0\), we do get \(y=\pm 6\)

ganeshie8 (ganeshie8):

that confirms the equation has correct vertices

OpenStudy (bluebeta):

WOW!! Thanks!! Did you look over the file I posted?? I just wanna make sure everything else is good :) Please and THANK YOU SO MUCH!!

ganeshie8 (ganeshie8):

#10 looks good

ganeshie8 (ganeshie8):

i haven't checked first 8 problems

OpenStudy (bluebeta):

Do you care too?

ganeshie8 (ganeshie8):

#1 is circle I think you need to find the radius too

ganeshie8 (ganeshie8):

#2, #3, #4, #5 look good

OpenStudy (bluebeta):

Do I need the radiisu for 1-4?

OpenStudy (bluebeta):

@ganeshie8